# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v1

@@ C.5 Value function approximation
Let $K^{\prime}\geq 0$ be a macro-iteration. Under the event that $|V_{\pi_{K^{\prime}-1}^{t},p_{K^{\prime}-1}}^{i}(\rho)-V_{\pi_{K^{\prime}-1}^{t},p_{K^{\prime}-1}}^{i}(\rho)|\leq\epsilon$ for all $i\in[m]$ and $\Delta_{p}=\Theta\left(\frac{(1-\gamma)}{\alpha\log(A)}\right)$, it follows that $\left\lVert\lambda_{K^{\prime}}\right\rVert_{1}=\mathcal{O}(1)$.
@@
The proof makes use of complementary slackness, telescoping, the inner product in Lemma 13, and the basic upper bound via property 3 of Lemma 12 that $\left\lVert\lambda_{0}\right\rVert\leq\left\lVert\lambda_{0}\right\rVert_{1}\leq\eta_{\lambda}^{2}\frac{m}{1-\gamma}$. Additionally, due to the slack variable $\zeta>0$ in Assumption 3 for any $p\in\mathcal{P}$ and Lemma 16 in Liu et al., (2021), the optimal dual variable can be bounded based on $\lambda^{*}\leq\frac{2}{(1-\gamma)\zeta}=\mathcal{O}(1)$. The full derivation is rather lengthy so we refer to Lemma 21 of Liu et al., (2021). Our analysis adds a correction for the transition dynamic changes within the uncertainty set $\Delta_{p}=\mathcal{O}(1)$ as derived in Eq. 43. Since the additional term is $\alpha\Delta_{p}\mathcal{O}(\frac{\log(A)}{1-\gamma})=\mathcal{O}(1)$, we obtain the desired result. ∎
Appendix D Algorithm implementation details
@@
All the algorithms were implemented in Pytorch. The code for RMCPMD was taken from its original implementation https://github.com/JerrisonWang/JMLR-DRPMD and the code for PPO was taken from the StableBaselines3 class. Both were modified to fit our purposes by allowing to turn off and on robust training and constraint-satisfaction. Similar to PPO, MDPO was also implemented following StableBaselines3 class conventions. All the environments are implemented in Gymnasium. The updates with LTMA are based on modifying the TMA code in https://github.com/JerrisonWang/JMLR-DRPMD. For MDPO and PPO experiments, we use four parallel environments and with four CPUs. Our longest experiments typically take no more than two hours to complete. The remainder of the section describes domain-specific hyperparameter settings.
D.1 Hyperparameters for Cartpole experiments
@@
D.2 Hyperparameters for Inventory Management experiments
Hyperparameters settings for the unidimensional Inventory Management (see Table 9) are similar to the Cartpole experiments. The key differences are the output layer, the discount factor, the transition kernel, and the TMA parameters, which follow settings of the domain in Wang and Petrik, (2024). Based on initial tuning experiments with PPO and MDPO, the learning rates are set to $\eta=1\mathrm{e}^{-3}$ and $\eta_{\lambda}=1\mathrm{e}^{-2}$, a lower GAE lambda of 0.50 was chosen and PPO is implemented without early stopping and with the same number of (5) epochs as MDPO.
@@ D.2 Hyperparameters for Inventory Management experiments
D.3 Hyperparameters for 3-D Inventory Management experiments
@@
Appendix E Training performance plots
@@ E.2 Inventory Management
Appendix F Test performance plots with large sample budget
@@ F.1 Cartpole
Figure 13: Test performance of MDP and CMDP algorithms obtained by applying the learned determinstic policy from the Cartpole domain after 200,000 time steps of training. The line and shaded area represent the mean and standard error across the perturbations for the particular distortion level.
@@
Figure 14: Test performance of RMDP and RCMDP algorithms obtained by applying the learned determinstic policy from the Cartpole domain after 200,000 time steps of training. The line and shaded area represent the mean and standard error across the perturbations for the particular distortion level.
F.2 Inventory Management
@@
Figure 15: Test performance of MDP and CMDP algorithms obtained by applying the learned determinstic policy from the Inventory Management domain after 400,000 time steps of training. The line and shaded area represent the mean and standard error across the perturbations for the particular distortion level.
@@
Figure 16: Test performance of RMDP and RCMDP algorithms obtained by applying the learned determinstic policy from the Inventory Management domain after 400,000 time steps of training. The line and shaded area represent the mean and standard error across the perturbations for the particular distortion level.
F.3 3-D Inventory Management
@@
Figure 17: Test performance of RMDP and RCMDP algorithms obtained by applying the learned determinstic policy from the 3-D Inventory Management domain after 400,000 time steps of training. The line and shaded area represent the mean and standard error across the perturbations for the particular distortion level.
@@
Figure 18: Test performance of RMDP and RCMDP algorithms obtained by applying the learned determinstic policy from the 3-D Inventory Management domain after 400,000 time steps of training. The line and shaded area represent the mean and standard error across the perturbations for the particular distortion level.
Appendix G Comparison of schedules
Table 11 summarises the experiments comparing restart schedules to standard linear schedules as well as geometric schedules for the $\alpha_{k}^{-1}$ parameter in MDPO-Robust-Lag variants.

### v2

@@ D.6.2 Main theorem
Let $K^{\prime}\geq 1$ be a macro-iteration. Under the event that $|V_{\pi_{K^{\prime}-1},p_{K^{\prime}-1}}^{i}(\rho)-V_{\pi_{K^{\prime}-1},p_{K^{\prime}-1}}^{i}(\rho)|\leq\epsilon$ for all $i\in[m]$, it follows that $\left\lVert\lambda_{K^{\prime}}\right\rVert_{1}=\mathcal{O}(1)$.
@@
Consider the Lagrangian with optimal parameters $(\pi^{*},\lambda^{*},p^{*})$. Note that
$K^{\prime}V_{\pi^{*},p^{*}}(\rho;\lambda^{*})$ $=K^{\prime}\mathbf{V}_{\pi^{*},p^{*}}(\rho;\lambda^{*})$ (complementary slackness)
$\leq\sum_{k=1}^{K^{\prime}}\mathbf{V}_{\pi_{k},p^{*}}(\rho;\lambda^{*})$
$=\sum_{k=1}^{K^{\prime}}V_{\pi_{k},p^{*}}(\rho)+\sum_{j=1}^{m}\lambda_{j}^{*}V_{\pi_{k},p^{*}}^{j}(\rho)$
$\leq\sum_{k=1}^{K^{\prime}}V_{\pi_{k},p_{k-1}^{*}}(\rho)+\sum_{j=1}^{m}\lambda_{j}^{*}V_{\pi_{k},p_{k-1}^{*}}^{j}(\rho)$ ($p_{k-1}^{*}$ maximises the Lagrangian for $\pi_{k}$)
$\leq\frac{M}{1-\gamma}\left(\sum_{k=1}^{K^{\prime}}V_{\pi_{k},p_{k}}(\rho)+\sum_{j=1}^{m}\lambda_{j}^{*}V_{\pi_{k},p_{k}}^{j}(\rho)\right)$ (Lemma 26)
$\leq\frac{M}{1-\gamma}\left(\mathcal{O}(1)+\sum_{k=1}^{K^{\prime}}V_{\pi_{k},p_{k}}(\rho)+\sum_{j=1}^{m}\lambda_{j}^{*}\left(\frac{1}{\eta_{\lambda}}\lambda_{K^{\prime},j}-\sum_{k=1}^{K^{\prime}}\epsilon_{k,j}\right)\right)$ (derivations in Eq. 74)
$\leq\frac{M}{1-\gamma}\left(\mathcal{O}(1)+\sum_{j=1}^{m}\lambda_{j}^{*}\frac{1}{\eta_{\lambda}}\lambda_{K^{\prime},j}-\sum_{k=1}^{K}\sum_{j=1}^{m}\lambda_{j}^{*}\epsilon_{k,j}\right$.
$+\sum_{k=1}^{K^{\prime}}\frac{\alpha}{1-\gamma}\left(B_{d_{\rho}^{\pi^{*},p_{k-1}}}(\pi^{*},\pi_{k-1})-B_{d_{\rho}^{\pi^{*},p_{k-1}}}(\pi^{*},\pi_{k})\right)+\Delta_{k}$
$\left.+\frac{\eta_{\lambda}}{2}\left(\left\lVert V_{\pi_{k},p_{k-1}}^{1:m}(\rho)\right\rVert^{2}-\left\lVert V_{\pi_{k-1},p_{k-1}}^{1:m}(\rho)\right\rVert^{2}\right)+\frac{1}{2\eta_{\lambda}}\left(\left\lVert\lambda_{k-1}\right\rVert^{2}-\left\lVert\lambda_{k}\right\rVert^{2}\right)\right)\,.$ (following Eq. 72)
Applying Lemma 25 to $p^{*}$, it follows that $\left\lVert\lambda^{*}\right\rVert=\mathcal{O}(1)$ such that $\sum_{j=1}^{m}\lambda_{j}^{*}\mathcal{O}(\epsilon)=\mathcal{O}(\epsilon)$. Note then that via Lemma 24, $\frac{\alpha}{1-\gamma}\sum_{k\in[K]}\left(B_{d_{\rho}^{\pi^{*},p_{k-1}}}(\pi^{*},\pi_{k-1})-B_{d_{\rho}^{\pi^{*},p_{k-1}}}(\pi^{*},\pi_{k})\right)\leq\frac{\alpha}{1-\gamma}\left(\log(A)+2\log(\frac{1}{U})\right)$. Moreover, the preconditions for $\alpha$ in the further derivations in Lemma 21 of Liu et al., 2021b are satisfied. The remainder of the proof removes the constant $\frac{M}{1-\gamma}=\mathcal{O}(1)$ and then follows the derivations in Lemma 21 of Liu et al., 2021b . ∎
The lemma below bounds the error induced by the dual variable taking suboptimal values during policy optimisation and LTMA. The result bounds the instantaneous additional error by a constant, and the average error as $\mathcal{O}(\epsilon)$.
Lemma 28 (Bound on dual variable induced error).
Let $k=1,\dots,K$ define a sequence of macro-iterations with the preconditions of Theorem 2. Let $(p_{k-1}^{*},\lambda_{k-1}^{*})$ be the optimal solutions for $\lambda$ and $p$ in the problem
$\max_{\lambda\geq 0}\sup_{p\in\mathcal{P}}\left(V_{\pi_{k},p}(\rho)+\sum_{j=1}^{m}\lambda_{j}V_{\pi_{k},p}^{j}(\rho)\right)\,.$
Then the error induced by the dual variable, as defined in Lemma 8 according to $\Delta_{\lambda_{k}}=\sum_{j=1}^{m}(\lambda_{k-1,j}^{*}-\lambda_{k-1,j})V_{\pi_{k},p_{k-1}^{*}}^{j}(\rho)$ satisfies
$\frac{1}{K}\sum_{k=1}^{K}\Delta_{\lambda_{k}}=\mathcal{O}(\epsilon)\,$ (75)
Proof.
Note that
$\frac{1}{K}\sum_{k=1}^{K}\Delta_{\lambda_{k}}$ $=\frac{1}{K}\sum_{k=1}^{K}\sum_{j=1}^{m}(\lambda_{k-1,j}^{*}-\lambda_{k-1,j})V_{\pi_{k},p_{k-1}^{*}}^{j}(\rho)$
$\leq\frac{1}{K}\sum_{k=1}^{K}\sum_{j=1}^{m}B_{\lambda}V_{\pi_{k},p_{k-1}^{*}}^{j}(\rho)$ (non-negativity in Lemma 13 and Assumption 2)
$\leq\frac{M}{1-\gamma}B_{\lambda}\frac{1}{K}\sum_{k=1}^{K}\sum_{j=1}^{m}V_{\pi_{k},p_{k}}^{j}(\rho)$ (Lemma 26)
$=\mathcal{O}(\epsilon)\,.$ (Theorem 2b))
Appendix E Algorithm implementation details
@@
All the algorithms were implemented in Pytorch. The code for RMCPMD was taken from its original implementation https://github.com/JerrisonWang/JMLR-DRPMD and the code for PPO was taken from the StableBaselines3 class. Both were modified to fit our purposes by allowing to turn off and on robust training and constraint-satisfaction. Similar to PPO, MDPO was also implemented following StableBaselines3 class conventions. All the environments are implemented in Gymnasium. For simplicity, the updates with LTMA are based on the TMA code in https://github.com/JerrisonWang/JMLR-DRPMD, which performs pure Monte Carlo based TMA rather than additional function approximation as proposed in our Approximate TMA algorithm. For MDPO and PPO experiments, we use four parallel environments and with four CPUs. Our longest experiments typically take no more than two hours to complete. The remainder of the section describes domain-specific hyperparameter settings.
E.1 Hyperparameters for Cartpole experiments
@@
E.2 Hyperparameters for Inventory Management experiments
Hyperparameters settings for the unidimensional Inventory Management (see Table 9) are similar to the Cartpole experiments. The key differences are the output layer, the discount factor, the transition kernel, and the TMA parameters, which follow settings of the domain in Wang et al., (2024). Based on initial tuning experiments with PPO and MDPO, the learning rates are set to $\eta=1\mathrm{e}^{-3}$ and $\eta_{\lambda}=1\mathrm{e}^{-2}$, a lower GAE lambda of 0.50 was chosen and PPO is implemented without early stopping and with the same number of (5) epochs as MDPO.
@@ E.2 Hyperparameters for Inventory Management experiments
E.3 Hyperparameters for 3-D Inventory Management experiments
@@
Appendix F Training performance plots
@@ F.2 Inventory Management
Appendix G Test performance plots with large sample budget
@@ G.1 Cartpole
Figure 13: Test performance of MDP and CMDP algorithms obtained by applying the learned deterministic policy from the Cartpole domain after 200,000 time steps of training. The line and shaded area represent the mean and standard error across the perturbations for the particular distortion level.
@@
Figure 14: Test performance of RMDP and RCMDP algorithms obtained by applying the learned deterministic policy from the Cartpole domain after 200,000 time steps of training. The line and shaded area represent the mean and standard error across the perturbations for the particular distortion level.
G.2 Inventory Management
@@
Figure 15: Test performance of MDP and CMDP algorithms obtained by applying the learned deterministic policy from the Inventory Management domain after 400,000 time steps of training. The line and shaded area represent the mean and standard error across the perturbations for the particular distortion level.
@@
Figure 16: Test performance of RMDP and RCMDP algorithms obtained by applying the learned deterministic policy from the Inventory Management domain after 400,000 time steps of training. The line and shaded area represent the mean and standard error across the perturbations for the particular distortion level.
G.3 3-D Inventory Management
@@
Figure 17: Test performance of RMDP and RCMDP algorithms obtained by applying the learned deterministic policy from the 3-D Inventory Management domain after 400,000 time steps of training. The line and shaded area represent the mean and standard error across the perturbations for the particular distortion level.
@@
Figure 18: Test performance of RMDP and RCMDP algorithms obtained by applying the learned deterministic policy from the 3-D Inventory Management domain after 400,000 time steps of training. The line and shaded area represent the mean and standard error across the perturbations for the particular distortion level.
Appendix H Comparison of schedules
Tomar et al., (2022) propose either a fixed $\alpha$ or a linear increase across iterations to give a larger penalty at the end of the iterations (inspired by theoretical works, e.g. Beck and Teboulle, (2003)). A potential problem with such a linear scheme is that if the LTMA update is strong and the policy cannot take sufficiently large learning steps, it may sometimes lead to a gradual loss of performance and premature convergence, especially in the context of contrained optimisation where the objective is often subject to large changes due to constraints becoming active or inactive. As an alternative, we also consider a schedule where the penalty parameter schedule is restarted every time the LTMA update to $\xi$ is larger than a particular threshold. The schedule then follows the process $\alpha_{k}=\frac{1}{1-(k-k^{\prime})/(K-k^{\prime})}$, where $k^{\prime}$ indicates the restart time and $K$ is the final time. In the experiments, the restart is done every time any dimension $i$ of the update is larger than half the maximal distance from nominal, i.e. larger than $0.5\max_{\xi_{i}\in\mathcal{U}_{\xi_{i}}}|\xi_{i}-\bar{\xi}_{i}|$. A last included schedule is the geometrically decreasing schedule $\alpha_{k}=\alpha_{k-1}*\gamma$ based on the discount factor $\gamma$ as proposed in earlier work (Xiao,, 2022). We experiment with both epoch-based and batch-based schedules. While in the batch-based schedules we additionally experimented with the more aggressive update rule $\alpha_{k}=\alpha_{k-1}\frac{1-\gamma}{M}$ proposed in Wang et al., (2024). However, the results under the scheme were poor, and so these results are not included.
Table 11 and 12 summarise the experiments comparing fixed schedules, restart schedules, linear schedules, and the geometric schedules for MDPO-Robust and MDPO-Robust-Lag variants. For the fixed schedule, the parameter is set to $\alpha=2.0$. For the geometric schedule, the starting parameter is set to $\alpha=5.0$ for epoch-based scheduling and to $\alpha=2.0$ for the less frequent batch-based scheduling. The fixed setting works reasonably well while in particular cases there may be benefits from time-varying schedules. The best overall performance for robust-constrained RL is obtained by the batch-based restart schedule with MDPO-Robust-Augmented-Lag. As a second top performer, the geometric schedule with MDPO-Robust-Lag also performs consistently high.